OpenStudy (anonymous):

simplify the square root of 20 puls the square foot of 40

OpenStudy (anonymous):

can anyone help not sure how to solve?

OpenStudy (unklerhaukus):

helo, welcome to openstudy!

OpenStudy (unklerhaukus):

\[\sqrt {20}+\sqrt{40}\] simplify the numbers by factoring \[20 = 4\times5=2^2\times5 \\40=\]

OpenStudy (unklerhaukus):

can you factor 40 into prime factors ?

OpenStudy (anonymous):

yes it would be 4*10=2^3*5 right?

OpenStudy (unklerhaukus):

Right.

OpenStudy (anonymous):

i am not really sure what there is to "simplify" here you can write \(\sqrt{20}=\sqrt{4\times 5}=\sqrt{4}\times \sqrt{5}=2\sqrt{5}\) which puts it in simplest radical form

OpenStudy (unklerhaukus):

so we can write\[\sqrt {20}+\sqrt{40}=\sqrt{2^2\times5}+\sqrt{2^3\times5}\] now if there are terms under a square root that are squared, we can take them out, and drop the square , \[=2\sqrt{5}+\sqrt{2^3\times5}\] for the second radical we need to break up the 2^3 so we can extract the square term , 2^3 = (some square term) times (something else)

OpenStudy (unklerhaukus):

can you simplify the second term @iun45 ?

OpenStudy (anonymous):

2^2*2 right

OpenStudy (unklerhaukus):

good,

OpenStudy (unklerhaukus):

now take the square term out,

OpenStudy (anonymous):

so we get 2

OpenStudy (unklerhaukus):

do you mean =2√5+2√(2×5) ?

OpenStudy (anonymous):

2^2√5+2^2√5*2 ?

OpenStudy (unklerhaukus):

the ^2 bit cancels with the √ bit when you bring the term out from under the racial so \[\quad\sqrt{2^2\times 5}+\sqrt{2^2\times 5\times2}\\=\sqrt{2^2}\times\sqrt{ 5}+\sqrt{2^2}\times \sqrt{5\times2}\\=2\times\sqrt{ 5}+2\times \sqrt{5\times2}\\=2\sqrt{ 5}+2 \sqrt{5\times2}\]

OpenStudy (unklerhaukus):

i have used this for simplifying the square roots\[\sqrt{ab}=\sqrt a\sqrt b\] use it again to factor the last term,

OpenStudy (unklerhaukus):

finally factor out the common factors of each term \[ab+abc=ab(1+c)\]

OpenStudy (anonymous):

ok now i'm lost